Assume the following:

- 20% of the hospital’s bills are paid on day 30

- 40% of the hospital’s bills are paid on day 60

- 28% of the hospital’s bills are paid on day 75

- The rest are paid on day 90

A) What is the hospital’s average collection period?_____________

B) If the hospital’s annual patient service revenue is $10,000,000 (assume 360 days), what is the current receivables balance? _____________________
Answer #1

A) The average collection Period =20%*30+40%*60+28%*75+12%*90=61.80

b) Current Receivables balance =Annual Revenue* Average Collection Period/365 =10000000*61.80/365 =1,693,150.68